OpenStudy (anonymous):

|7/5-x/5|=1/5 Help pleaseee xD

OpenStudy (anonymous):

\[|\frac{7}{5}-\frac{x}{5}|=\frac{1}{5}\] \[|\frac{1}{5}(7-x)|=\frac{1}{5}\] \[\frac{1}{5}|7-x|=\frac{1}{5}\] \[|7-x|=1\] \[7-x=1\] or \[7-x=-1\] \[x=6\] or \[x=8\]

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

yw